OpenStudy (anonymous):

Seriously need help!!(3^-1x^-1y^-1)^2(3x^-4y^3)(9x^-2y^3)0/(3x^-3y^-5)^2 simplify, step by step help!!!

OpenStudy (aum):

\[ \frac{(3^{-1}x^{-1}y^{-1})^2(3x^{-4}y^3)(9x^{-2}2y^3)^0}{(3x^{-3}y^{-5})^2} \]Is that the correct problem? Is the third factor in the numerator raised to zero?

OpenStudy (anonymous):

yes and yes.

OpenStudy (aum):

\[ \frac{(3^{-1}x^{-1}y^{-1})^2(3x^{-4}y^3)(9x^{-2}2y^3)^0}{(3x^{-3}y^{-5})^2} = \frac{(3^{-2}x^{-2}y^{-2})*(3x^{-4}y^3)*1}{(9x^{-6}y^{-10})} = \\ \frac{3^{-2+1}x^{-2-4}y^{-2+3}}{9x^{-6}{-10}} = ? \]Can you simplify from here on?

OpenStudy (anonymous):

yes thank you very much!

OpenStudy (aum):

you are welcome.

OpenStudy (aum):

\[ \Large \frac{3^{-2+1}x^{-2-4}y^{-2+3}}{9x^{-6}{y^{-10}}} = ? \]

OpenStudy (anonymous):

it would be y/3(9-10x^6)? I think

OpenStudy (aum):

\(\Large \frac{3^{-2+1}x^{-2-4}y^{-2+3}}{9x^{-6}{y^{-10}}} = \frac{3^{-1}x^{-6}y^{1}}{9x^{-6}{y^{-10}}} = \frac{y^{1}}{3*9*y^{-10} } = \frac{y^{1+10}}{27} = \frac{y^{11}}{27} \)

OpenStudy (anonymous):

Oh alright I get thank you.